What is the standard formula for the quadratic function that has a vertex of (0.5,4) and passes through points (-1.5,0) and (2.5,0)?
Answer by lwsshak3(11628)

What is the standard formula for the quadratic function that has a vertex of (0.5,4) and passes through points (-1.5,0) and (2.5,0)?
**
Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of vertex.
For given function:
y=A(x-0.5)^2+4
solve for A using one of given points (-1.5,0)
0=A(-1.5-0.5)^2+4
0=A(-2)^2+4
4A=-4
A=-1
Equation: y=-(x-0.5)^2+4